Theresa has $1.75 in quarters and nickles. If she has 15 coins total, how many quarters does Theresa have?
Thank you!

Answer by stanbon(75887)

Theresa has $1.75 in quarters and nickles. If she has 15 coins total, how many quarters does Theresa have?
--------------
Quantity Eq: n + q = 15 coins
Value Eq: 5n+25q = 175 cents
--------------------------------
Divide thru the Value Eq. by 5.
---
n + q = 15
n +5q = 35
----------------
Subtract and solve for "q":
4q = 20
q = 5 (# of quarters)
------------------------
Solve for "n":
n + q = 15
n + 5 = 15
n = 10 (# of nickels)
===========================
Cheers,
Stan H.
===========================